OpenStudy (anonymous):

I have an aluminum cylinder; length =10cm and radius =.25cm density of aluminum =2.70g/cm3 1Al atom =4.48*10-23g How many Al Atoms are in the cylinder? Would you find the volume first? then do mass = d*v? and divide that by the mass of 1 Al atom?

OpenStudy (anonymous):

@toxicsugar22 do you know?

OpenStudy (anonymous):

@3-sfera doing what i explained in the question i got 1.18*10^23 atoms of Al, do you know if i did this correctly

OpenStudy (anonymous):

let me do some math

OpenStudy (anonymous):

no my result is 1,2x10^27

OpenStudy (anonymous):

i did Vcylinder = pi r^2(h) pi(.25^2)(10) = 1.96cm^3 then mass = d*v m= 2.70*1.96 = 5.292 then 5.292 / 4.48*10^-23 = 1.18*10^23

OpenStudy (anonymous):

in the question it says taht 1 atom of Al = 4.48 * 10^-23 g

OpenStudy (anonymous):

my solution is with the molar mass, your volume is incorrect i think. the volume is 1,96*10^3 cm^3

OpenStudy (anonymous):

so my number of atoms is correct too? 1.18*10^23

OpenStudy (anonymous):

uhm actually not :).. maybe just some different approximation..maybe the mass on your book is wromg

OpenStudy (anonymous):

Well by the book, the answer is correct then?

OpenStudy (anonymous):

@aaronq can you please help?

OpenStudy (anonymous):

the answer with your atom mass is 1.8*10^26

OpenStudy (anonymous):

sorry for late i'm a little bit busy

OpenStudy (anonymous):

i don't understand how you got that.

OpenStudy (anonymous):

oh flutter wait.. the radius is 0.25

OpenStudy (anonymous):

*flutter

OpenStudy (anonymous):

ahha i wanna write f*u*c*k but it shows flutter

OpenStudy (anonymous):

by the way, i misread the radius

OpenStudy (anonymous):

your answer is correct i think

OpenStudy (anonymous):

sorry @Savannah_Lynn13